Stan's, Mark's, and Wayne's ages are consecutive whole numbers. Stan is the youngest, and Wayne is the oldest. The sum of their ages is 111. Find their ages.
Answer by richwmiller(17219)

1+2=3
111-3=108
108/3=36
36,37,38